OpenStudy (anonymous):

I don't know where to begin with this, -3+40sinx=32cos^2 x

OpenStudy (anonymous):

what about it

OpenStudy (anonymous):

@petewe I don't know how to solve this.

OpenStudy (anonymous):

-3 + 40sinx - 32(1-sin^2(x)) = 0 -3 + 40sinx - 32 + sin^2(x) = 0 40sinx + sin^2(x) -35 = 0 solve the quadratic.